OpenStudy (hayleymeyer):

What is the temperature, in degrees Celsius, of a substance with a temperature of 49K? –322°C –224°C 224°C 322°C

OpenStudy (hayleymeyer):

please help

OpenStudy (nayef):

Is this a homework or assignment?

OpenStudy (hayleymeyer):

an assignment, im having trouble with science

OpenStudy (nayef):

Use the equation: degree in Kelvin=degree in Celsius+273

OpenStudy (shamim):

U know 0 degree celcius=0+273=273k

OpenStudy (shamim):

1 degree celcius=1+273=274k

OpenStudy (shamim):

2 degree celcius=2+273=275k